\begin{thebibliography}{41}

\bibitem{Jiahui1999} Jiahui W 1999 Phase Structure of Z 3 Lattice Gauge Theory Coupled With the Higgs Field \textit{Chinese Physics C} \textbf{23} 352--359
\bibitem{Anber2013} Anber M M, Collier S and Poppitz E 2013 The SU (3)/Z3 QCD (adj) deconfinement transition via the gauge theory"affine" XY-model duality \textit{Journal of High Energy Physics} \textbf{2013} 1--26
\bibitem{Robaina2021} Robaina D , Ba{\~n}uls M C and Cirac J I 2021 Simulating 2+ 1 D Z 3 lattice gauge theory with an infinite projected entangled-pair state \textit{Physical Review Letters} \textbf{126} 050401
\bibitem{Baranka2022} Baranka G and Giordano M 2022 Deconfinement transition and localization of Dirac modes in finite-temperature Z 3 gauge theory on the lattice \textit{Physical Review D} \textbf{106} 094508
\bibitem{Akiyama2023} Akiyama S and Kuramashi Y 2023 Critical endpoint of (3+ 1)-dimensional finite density $ Z\_3 $ gauge-Higgs model with tensor renormalization group \textit{arXiv preprint arXiv:2304.07934} 
\bibitem{Celik1985} Celik T , G{\"u}nd{\"u}{\c{c}} Y and Satz H 1985 Finite-temperature deconfinement transition in (2+ 1)-dimensional Z 3 gauge theory \textit{Physical Review D} \textbf{32} 2246
\bibitem{Ko2015} Ko P and Tang Y 2015 Galactic center $\gamma$-ray excess in hidden sector DM models with dark gauge symmetries: local Z3 symmetry as an example \textit{Journal of Cosmology and Astroparticle Physics} \textbf{2015} 023
\bibitem{Choi2015} Choi S and Lee H M 2015 SIMP dark matter with gauged Z 3 symmetry \textit{Journal of High Energy Physics} \textbf{2015} 1--21
\bibitem{Choi2020} Choi S , Kim J , Lee H M and Zhu B 2020 Connecting between inflation and dark matter in models with gauged Z3 symmetry \textit{Journal of High Energy Physics} \textbf{2020} 1--39
\bibitem{Borah2023} Borah D , Ma E and Nanda D 2023 Dark SU(2) $\rightarrow$ Z3 $\times$ Z2 gauge symmetry \textit{Physics Letters B} \textbf{842} 137981
\bibitem{Drouffe1981} Drouffe J and Zuber J 1981 Roughening transition in lattice gauge theories in arbitrary dimension:(II). The groups Z3, U (1), SU (2), SU (3) \textit{Nuclear Physics B} \textbf{180} 264--274
\bibitem{Emonts2020} Emonts P , Ba{\~n}uls M C, Cirac I and Zohar E 2020 Variational Monte Carlo simulation with tensor networks of a pure Z 3 gauge theory in (2+ 1) D \textit{Physical Review D} \textbf{102} 074501
\bibitem{Gattringer2012} Gattringer C and Schmidt A 2012 Gauge and matter fields as surfaces and loops: An exploratory lattice study of the Z 3 gauge-Higgs model \textit{Physical Review D} \textbf{86} 094506-
\bibitem{Fisch1994} Fisch R 1994 Freezing transition in the Z3 gauge glass \textit{Journal of Applied Physics} \textbf{75} 5544--5546
\bibitem{Brower1991} Brower R C and Huang, S 1991 Dynamic universality for Z 2 and Z 3 lattice gauge theories at finite temperature \textit{Physical Review D} \textbf{44} 3911
\bibitem{Hattori2010} Hattori C , Matsunaga M , Matsuoka T and Nakanishi K 2010 $SU(3)_L \times (Z3 \times  Z3$ gauge symmetry and tri-bimaximal mixing \textit{Modern Physics Letters A} \textbf{25} 2981--2989
\bibitem{Brydges1982} Brydges D, Fr\"{o}hlich J and Spencer T 1982 The random walk representation of classical spin systems and correlation inequalities \textit{Communications In Mathematical Physics} \textbf{83} 123-150
\bibitem{Siqueira1986} Siqueira A and Fittipaldi I 1986 New effective-field theory for the Blume-Capel model \textit{Physica A: Statistical Mechanics and its Applications} \textbf{138} 592--611
\bibitem{SaBarreto2012} Barreto F S and Mota A 2012 Correlations equalities and some upper bounds for the critical temperature for spin one systems \textit{Physica A: Statistical Mechanics and its Applications} \textbf{391} 5908--5917
\bibitem{SaBarreto2015} Barreto F S and Mota A 2015 Area decay of Wilson loop for lattice gauge theories with bond variables described by the transverse Ising model \textit{Physica A: Statistical Mechanics and its Applications} \textbf{434} 144--150
\bibitem{Griffiths1969} Griffiths R B 1969 Rigorous results for Ising ferromagnets of arbitrary spin \textit{Journal of Mathematical Physics} \textbf{10} 1559--1565
\bibitem{Simon1980} Simon B, Pollard N, Srivastava R and Zakeri M 1980 Correlation inequalities and the decay of correlations in ferromagnets \textit{Communications In Mathematical Physics} \textbf{77} 111-126
\bibitem{Lieb1980} Lieb, E H 1980 A refinement of Simon's correlation inequality \textit{Communications in Mathematical Physics} \textbf{77} 127--135.
\bibitem{Ginibre1969} Ginibre J 1969 Simple proof and generalization of Griffiths' second inequality \textit{Phys. Rev. Lett.} \textbf{23} 828--830
\bibitem{Sylvester1976} Sylvester G S 1976 Inequalities for continuous-spin Ising ferromagnets \textit{Journal of Statistical Physics} \textbf{15} 327--341
\bibitem{Ganikhodjaev2007} Ganikhodjaev N and Razak F A 2007 Correlation Inequalities for Generalized Potts Model: General Griffiths' Inequalities \textit{arXiv preprint arXiv:0707.3848}
\bibitem{Ganikhodjaev2010} Ganikhodjaev N and Razak F A 2010 Griffith--Kelly--Sherman correlation inequalities for generalized Potts model \textit{Mathematical Physics, Analysis and Geometry} \textbf{13} 1--18
\bibitem{Newman1975} Newman C M 1975 Gaussian correlation inequalities for ferromagnets \textit{Zeitschrift f{\"u}r Wahrscheinlichkeitstheorie und Verwandte Gebiete} \textbf{33} 75--93
\bibitem{Wilson1974} Wilson K G 1974 Confinement of quarks \textit{Physical review D} \textbf{10} 2445
\bibitem{Seiler1982} Seiler E 1982 Gauge theories as a problem of constructive quantum field theory and statistical mechanics \textit{Lect. Notes Phys.} \textbf{159} 1--192
\bibitem{Kogut1979} Kogut J B 1979 An introduction to lattice gauge theory and spin systems \textit{Reviews of Modern Physics} \textbf{51} 659
\bibitem{Szasz1978} Sz\`{a}sz D 1978 Correlation inequalities for non-purely-ferromagnetic systems \textit{Journal Of Statistical Physics} \textbf{19} 453-459
\bibitem{SaBarreto1983a} Barreto F S and O'Carroll M 1983 Correlation equalities and coupling constant bounds implying area decay of Wilson loop for Z2 lattice gauge theories \textit{Journal of Physics A: Mathematical and General} \textbf{16} L431
\bibitem{Glimm1981} Glimm J , Jaffe A , Glimm J and Jaffe A 1981 Quantum Theory \textit{Quantum Physics: A Functional Integral Point of View} \textbf{} 3--27
\bibitem{SaBarreto1983b} Barreto F S and O'Carroll M 1983 Correlation equalities and some upper bounds for the critical temperature of Ising spin systems \textit{Journal of Physics A: Mathematical and General} \textbf{16} 1035
\bibitem{Tomboulis1981} Tomboulis E , Ukawa A and Windey P 1981 Correlation inequalities and quark confinement in lattice gauge theories \textit{Nuclear Physics B} \textbf{180} 294--300
\bibitem{Griffiths1967a} Griffiths R B 1967 Correlations in Ising ferromagnets. I \textit{Journal of Mathematical Physics} \textbf{8} 478--483
\bibitem{Griffiths1967b} Griffiths R B 1967 Correlations in Ising ferromagnets. II. External magnetic fields \textit{Journal of Mathematical Physics} \textbf{8} 484--489
\bibitem{Griffiths1967c} Griffiths R B 1967 Correlations in Ising ferromagnets. III: A mean-field bound for binary correlations \textit{Communications in Mathematical Physics} \textbf{6} 121--127
\bibitem{Braga1993} Braga G A, Ferreira S and Barreto F S 1993 Correlation Inequalities and Upper Bounds on the Critical Temperature for Some Spin Models \textit{Brazilian Journal of Physics} \textbf{23} 343-355
\bibitem{Braga1994} Braga G A, Ferreira S and Barreto F S 1994 Upper bounds on the critical temperature for the two-dimensional Blume-Emery-Griffiths model \textit{Journal of statistical physics} \textbf{76} 819--834
\end{thebibliography}

\documentclass[aps,nofootinbib]{revtex4-2}
\usepackage{epsfig}
\usepackage{epsf}

\begin{document}

\title{Correlations equalities and  some upper bounds for the coupling constant  implying area decay of Wilson loop for $Z_3$ lattice gauge theories}

\author{A. L. Mota}
\email{motaal@ufsj.edu.br}
\affiliation{Departamento de Ci\^{e}ncias Naturais, Universidade Federal de S\~{a}o Jo\~{a}o del Rei, \\
C.P. 110,  CEP 36301-160, S\~ao Jo\~ao del Rei, Brazil}
\author{F. C. S\'{a} Barreto}
\email{fcsabarreto@gmail.com}
\affiliation{Departamento de F\'{i}sica, Universidade Federal de Minas Gerais, C.P. 110, CEP 31270-901, Belo Horizonte, MG, Brazil}

\begin{abstract}
Correlation identities are obtained for $Z_3$ lattice gauge theory where the bonds of the plaquettes are decorated by generalized three-state Ising variables. Making use of correlation inequalities we obtain the area decay of the Wilson loop observable in a range of the coupling parameter larger than those obtained from mean field theory considerations.
\end{abstract}

\keywords{Correlation equalities, Rigorous upper bounds, $Z_3$ lattice gauge theory}
\maketitle

\section{Introduction}

Correlation inequalities combined with exact identities are useful in obtaining rigorous results in statistical mechanics. Among the various questions that are resolved by them one is the decay of the correlation functions. The decay of the correlation functions give information about the critical couplings of the statistical mechanics model. In this publication, the method will be applied to study the $Z_3$ lattice gauge problem where the bond variables are decorated by generalized three-state Ising variables $e^{n 2\pi i/3}$, n=0,1,2. The $Z_3$ discrete gauge symmetry is relevant to many fields of physics, such as, for instance, in the study of confinement/deconfinement transition in particle physics \cite{Jiahui1999,Anber2013,Robaina2021,Baranka2022,Akiyama2023,Celik1985}, in dark matter models, where the $Z_3$ symmetry ensures stability of dark matter particles \cite{Ko2015,Choi2015,Choi2020,Borah2023}, and, of course, in the study of lattice gauge theories and their transitions \cite{Drouffe1981,Emonts2020,Gattringer2012,Fisch1994,Brower1991,Hattori2010}. 

In this work, we firstly obtain correlation identities for the  $Z_3$ gauge lattice, which are the gauge version of the spin identities obtained previously for the spin one Blume-Capel model (\cite{Brydges1982,Siqueira1986,SaBarreto2012,SaBarreto2015}). We then look for the area decay of Wilson's loop, which is a criterion for confinement. We apply  the first and second Griffiths inequalities \cite{Griffiths1969,Ginibre1969,Sylvester1976,Simon1980,Ganikhodjaev2007,Ganikhodjaev2010} and Newman's inequalities \cite{Newman1975, Sylvester1976} to the correlation identities to obtain the coupling parameter which gives the area decay of the Wilson loop in d=3 and d=4 dimensions. In both dimensions the coupling parameters obtained for the area decay of the Wilson loop improve over the mean field results.

We consider the well known Wilson loop observable of a $ Z_3 $  pure gauge $ Z^d $ lattice theory with Wilson action defined as
\begin{equation}
\langle W( C )\rangle  = \lim_{\Lambda \to Z^d} \langle  W( C) \rangle _\Lambda \label{eq1}  
\end{equation}               
where $\langle W(C)\rangle $ is the finite lattice Gibbs ensemble average.
The Wilson action Boltzmann factor \cite{Wilson1974,Seiler1982,Kogut1979},is given by,
\begin{equation}
e^{{\beta J\sum_{P\subset\Lambda}\chi_P}  } \label{eq2}
\end{equation}                
where P denotes the unit squares (plaquettes) of $\Lambda$ and $\chi _P$ = $S_1S_2S_3S_4$. 
We let $S_b$ denote the bond variables which take values $e^{0}=1, e^{2\pi i/3}, e^{4\pi i/3}$ . W(C) is the product of $S_b$ along the perimeter of the planar rectangle C of area A. The gauge coupling constant is $\beta  J$ , $0 <  \beta  J < \infty$. 
Area decay of $\langle W( C )\rangle $ is a criterion for confinement. For the $Z_2$ lattice gauge theory it is known that area decay occurs for small $\beta J$ and for sufficiently large $\beta J$ perimeter decay occurs for $d \geq  3$ \cite{Seiler1982, Kogut1979}. For $d = 2$, $\langle W( C )\rangle $  has area decay for all $\beta J$ by explicit calculation \cite{Kogut1979}. We take free boundary conditions and note that Griffiths first and second inequalities  apply (\cite{Szasz1978, Griffiths1967a,Griffiths1967b,Griffiths1967c}), and especially for the Potts model see (\cite{Ganikhodjaev2007,Ganikhodjaev2010}), and therefore imply the existence of the thermodynamic limit \cite{Glimm1981}.
For d = 3 and 4 we obtain lower bounds $\beta_L$  on the area decay of $\langle W( C )\rangle $ i,.e. for all $\beta < \beta_L$, $\langle W ( C )\rangle $ has area decay, using correlation identities and Griffiths inequalities.  The correlation identities are a gauge version of the identities  presented before \cite{Siqueira1986,SaBarreto2012,SaBarreto2015} and employed to obtain lower than mean field upper bounds on the critical temperature for the Blume - Capel spin 1 systems. The procedure is similar to that employed previously \cite{SaBarreto1983a} to obtain the $Z_2$ lattice gauge correlation identities as a version of Callen's identities for the Ising spin 1/2 systems. 
For completeness, in section \ref{section2}, we give the mean field lower bound $\beta _M$ $(\beta_M < \beta_L)$ using a decoupling and Griffiths inequality argument \cite{SaBarreto1983b, Tomboulis1981}.

\section{Mean field lower bounds} \label{section2}
For definiteness assume that C lies in the ($x_1$, $x_2$)  plane. Consider a bond b fixed in the lower left-hand corner of C. Replace $\beta$  by $\beta \lambda$, $\lambda  \epsilon  [0,1]$, in the action for the $2(d-1)$ plaquettes (call them $P_1,P_2,...,P_{2(d-1)})$ that have one bond in common with b. Denote the corresponding expectation by $\langle W(C)\rangle _\lambda$, and let $J=1$, without loss of generality. Integrating $(d/d\lambda ) \langle W(C)\rangle _\lambda $ gives, noting that $\langle W(C)\rangle _0 = 0$,
\begin{eqnarray}
&& \langle W(C)\rangle  =\int_{0}^{1} d\lambda \frac{d}{d\lambda } \langle W(C)\rangle _\lambda = \beta \int_{0}^{1} d\lambda \sum_{i=1}^{2(d-1)} \left (\langle W(C) \chi _{P_i}\rangle _\lambda - \langle W(C)\rangle _\lambda \langle  \chi _{P_i}\rangle _\lambda \right)\nonumber \\
&&\leq \beta  \int_{0}^{1} d\lambda \sum_{i=1}^{2(d-1)} \langle W(C) \chi _{P_i}\rangle _\lambda \leq \beta \sum_{i=1}^{2(d-1)} \langle W(C) \chi _{P_i}\rangle ,\label{eq3}            
\end{eqnarray}             
In the previous we used Griffith's first (second) inequality in the first (second) equality. Each term on the right corresponds to a modified contour determined by the bonds of the variables of  $W( C ) \chi_{P_i}$ which enlarges or diminishes C by one plaquette. We repeat the argument proceeding along successive rows of plaquettes enclosed by C. After A applications we arrive at
\begin{equation}
\langle W( C )\rangle  \leq \beta^A \text{(sum of } [2(d-1)]^A \text{ terms)}         \label{eq4}   
\end{equation} 
Each term is non-negative and bounded above by 1 giving $\langle W( C )\rangle  \leq  [\beta  2(d-1)]^A$.
Therefore, for each $\beta$, 
\begin{eqnarray}
&&\beta \subset (0,(2(d-1)) ^ -1),\nonumber\\
&& \langle W( C )\rangle  \leq  e^{-\mid \ln (2\beta (d-1))\mid A}.     \label{eq5}  
\end{eqnarray} 

\section{ Correlation Identities for the $Z_3$ Lattice Gauge Model }

We define the action for the $Z_3$  lattice gauge model as
\begin{equation}
-H = \beta J\sum_{P\subset \Lambda }\chi _P  \label{eq21}
\end{equation}
where, $\chi _P$ is the square plaquette and $S_i$ is the bond variable assuming values  $e^{0}=1, e^{2\pi i/3} \equiv \sigma$, and $e^{4\pi i/3} = \sigma^2$.
Let $S_D = S_{i_1}...S_{i_D}$ denote a product of distinct bond variables and for a fixed bond b occurring in $S_D$ give a numerical ordering 1,2,... to the 2(d- 1) plaquettes that have one bond in common with b. 
We have,
\begin{equation}
\Big\langle S_D\Big\rangle =\frac{1}{Z} \sum_{S}S_D e^{\beta J\sum_{P\subset \Lambda }\chi _P }
\end{equation}
where,
\begin{equation}
Z= \sum_{{S}} e^{\beta J\sum_{P\subset \Lambda }\chi _P }
\end{equation}
Let us consider the bond b, with $S_b$, and the plaquette $\chi _b$ which contains the bond b.
Let $S_D^{(b)}$ be the product $S_D$ with the bond b deleted. We have,
\begin{equation}
\Big\langle S_D\Big\rangle =\frac{1}{Z}\cdot  \Big[\sum_{S}S_D^{(b)}\Big(\frac{\sum_{S_b}S_b e^{\beta J\chi _b}}{\sum_{S_b}e^{\beta J\chi _b }}\Big)\cdot e^{\beta J\sum_{P\subset \Lambda }\chi _P }\Big]\label{eq22}
\end{equation}
Let
\begin{equation}
\chi _b = S_b\sum_{neig.(b)}S_kS_lS_m
\end{equation}
Summing over $S_b$, in (\ref{eq22}) and introducing $\nabla$ = $\frac{\partial}{\partial x}$ through $e^ {a\nabla } f(x)=f(x+a)$, we get
\begin{equation}
\left\langle S_D\right\rangle =\left\langle S_D^{(b)}\left(\frac{e^{\alpha_b}+2 e^{-\frac{1}{2}\alpha_b} \cos\left( \frac{\sqrt{3}}{2} \alpha_b + \frac{2 \pi}{3} \right)}{e^{\alpha_b}+2 e^{-\frac{1}{2}\alpha_b} \cos\left( \frac{\sqrt{3}}{2} \alpha_b \right)}\right)\right\rangle
 = \left\langle S_D^{(b)}\prod_{neig(b)}e^{(\beta J S_k S_l S_m) \nabla} \right\rangle \cdot f(x)\mid_{x=0} \label{eq24} 
\end{equation}
where,  
\begin{equation}
\alpha_b = \beta J\sum_{neig.(b)} S_k S_l S_m
\end{equation}
and
\begin{equation}
f(x) = \frac{e^x + \sigma e^{\sigma x} + \sigma^2 e^{\sigma^2 x}}{e^x + e^{\sigma x} + e^{\sigma^2 x}} = \frac{e^x+2e^{-\frac{1}{2}x} \cos\left( \frac{\sqrt{3}}{2} x + \frac{2\pi}{3} \right)}{e^x+2e^{-\frac{1}{2}x} \cos\left( \frac{\sqrt{3}}{2} x \right)}.  \label{eqfx}
\end{equation}

Expanding the exponential in (\ref{eq24}) and using $S_i^{3n} = S_i^{3}=1$, $S_i^{3n+1} = S_i$, and $S_i^{3n+2} = S^2_i$, for $n=1,2,...$, we obtain,
\begin{equation}
\Big\langle S_D\Big\rangle =\Big\langle S_D^{(b)}\prod_{neig(b)}\Big[c_0(\beta J\nabla) + S_k S_l S_m s_1(\beta J\nabla ) + S^2_k S^2_l S^2_m s_2(\beta J\nabla ) \Big]
\Big\rangle \cdot f(x)\mid_{x=0}, \label{eq26} 
\end{equation}
with
\begin{equation}
c_0(x) = \frac{e^x+e^{\sigma x}+e^{\sigma^2 x}}{3}; \textrm{ }s_1(x) = \frac{e^x+\sigma^2 e^{\sigma x}+\sigma e^{\sigma^2 x}}{3}; \textrm{ and } s_2(x) = \frac{e^x+\sigma e^{\sigma x}+\sigma^2 e^{\sigma^2 x}}{3}. \label{c0s1s2}
\end{equation}

The functions $c_0(x), s_1(x)$, and $s_2(x)$ were named by their similarity to the hyperbolic $\cosh(x)$ and $\sinh(x)$ that appear in the equivalent treatment of the $Z_2$ lattice gauge \cite{SaBarreto1983a}. The following properties, derived from Eqs. (\ref{eqfx}) and (\ref{c0s1s2}) , are usefull in the evaluation of $\langle S_D \rangle$ in the next section:
\begin{equation}
f(\sigma x) = \sigma^2 f(x), \textrm{ } f(\sigma^2 x) = \sigma f(x),
\end{equation}
\begin{equation}
1 + \sigma + \sigma^2 = 0, \textrm{ } \sigma^3 = 1,
\end{equation}
and 
\begin{equation}
c^a_0(\beta J\nabla) s^b_1(\beta J\nabla) s^c_2(\beta J\nabla) f(x) \mid_{x=0} \Big\{ \begin{array}{c}\neq 0 \textrm{ if } (b+2c+1) = 3n  \\ = 0 \textrm{ otherwise},  \end{array}
\end{equation}
for $a, b, c,$ and $n$ positive integers.

In contrast with the $Z_2$ case \cite{SaBarreto1983a}, where it is given by the analytical function $tanh(x)$, $f(x)$ (Eq. (\ref{eqfx})) is non-analytical for $e^x+2e^{-\frac{1}{2}x} \cos\left( \frac{\sqrt{3}}{2} x \right) = 0$, i.e., at $x \approx -1.869812799$. However, as we shall see, this point is above both $d=3$ and $d=4$ upper limits for $\beta J$.

\section{ Exact correlation identities applied to the $d=3$, and $d=4$ lattices (square plaquettes)}

Applying Eq. (\ref{eq26}), for $d=3$ and $d=4$, after some algebra, we obtain, in subsections (\ref{ssd3}) and (\ref{ssd4}), with $k = \beta J$, $\rho_2 = 2 + \sigma$, $\rho_3 = 3 + \sigma$, $\rho_5 = 5 + \sigma$, and $\Re\{z\}$ stands for the real part of the complex argument $z$:

\subsection{d=3}\label{ssd3}
\begin{eqnarray}
\langle S_D\rangle &=&
A_1 \sum_{i} \langle S_D \chi^2_i \rangle 
+ A_2 \sum_{i<j} \langle S_D \chi_i \chi_j \rangle 
+ A_3 \sum_{i<j<k} \langle S_D \chi_i \chi^2_j \chi^2_k \rangle \nonumber \\
&+& A_4 \sum_{i<j<k<l} \langle S_D \chi_i \chi_j \chi_k \chi^2_l \rangle  
+ A_5 \sum_{i<j<k<l}\langle S_D \chi^2_i \chi^2_j \chi^2_k \chi^2_l\rangle, \label{SD3d}
\end{eqnarray}

with 

\begin{equation}
A_1=\frac{4}{9} \left( \frac{1}{3} f(4k) + f(k) - f(-2k) + 2\Re\{\rho_3 f(\rho_3 k)\} \right) \geq 0, \label{A1d3}
\end{equation}
\begin{equation}
A_2=\frac{2}{3} \left( \frac{1}{3} f(4k) + f(-2k) - \frac{4}{3} \Re\{\sigma f(\rho_3 k)\} \right) \geq 0,
\end{equation}
\begin{equation}
A_3=\frac{4}{3} \left( \frac{1}{3} f(4k) - f(k) + \frac{2}{3} \Re\{\sigma f(\rho_3 k)\} \right) \leq 0,
\end{equation}
\begin{equation}
A_4=\frac{4}{9} \left( \frac{1}{3} f(4k) + f(k) - f(-2k) + \frac{2}{3} \Re\{\sigma \rho_3 f(\rho_3^{*} k)\} \right) \leq 0,
\end{equation}
\begin{equation}
A_5=\frac{1}{9} \left( \frac{1}{3} f(4k) + 4 f(k) + 2 f(-2k) + \frac{8}{3} \Re\{\sigma f(\rho_3 k)\} \right) \geq 0. \label{A5d3}
\end{equation}

All inequalities hold for $0 \le k \le 0.924906$, where $f(-2k)$ is analytical, except Eq. (\ref{A1d3}), that holds for $0 \le k \le 0.740736$. Above this point we get $A_1 < 0$. As we shall show, the upper bound for the coupling constant lies in the interval of validity of Eqs. (\ref{A1d3})-(\ref{A5d3}).

\subsection{d=4}\label{ssd4}

\begin{eqnarray}
\langle S_D \rangle &=&
A_1 \sum_i \langle S_D \chi_i^2    \rangle + A_2 \sum_{i < j} \langle S_D \chi_i \chi_j   \rangle + A_3 \sum_{i < j<k} \langle S_D \chi_i \chi_j^2 \chi_k^2  \rangle \nonumber \\
&+& A_4 \sum_{i < j<k<l} \langle S_D \chi_i \chi_j \chi_k \chi_l^2   \rangle + A_5 \sum_{i<j<k<l<m} \langle S_D \chi_i \chi_j \chi_k \chi_l \chi_m   \rangle + A_6 \sum_{i<j<k<l} \langle S_D \chi_i^2 \chi_j^2 \chi_k^2 \chi_l^2   \rangle \nonumber \\ 
&+& A_7 \sum_{i<j<k<l<m} \langle S_D \chi_i \chi_j \chi_k^2 \chi_l^2 \chi_m^2   \rangle + A_8 \sum_{i < j<k<l<m<n} \langle S_D \chi_i \chi_j \chi_k \chi_l  \chi_m^2 \chi_n^2 \rangle \nonumber \\
&+& A_9 \sum_{i < j<k<l<m<n} \langle S_D \chi_i \chi_j^2 \chi_k^2 \chi_l^2  \chi_m^2 \chi_n^2 \rangle, \label{SD4d}
\end{eqnarray}
where the A coefficients are given by

\begin{equation}
A_1=\frac{2}{81} \left( f(6 k) + 15 f(3 k) - 10 f(-3 k) + 2 \Re \{ 10(2+\sigma) f(\rho_2 k) + (5+\sigma) f(\rho_5 k) + 5(2+\sigma) f(2 \rho_2 k) \}  \right) \ge 0, \label{A1d4}
\end{equation}

\begin{equation}
A_2=\frac{5}{81} \left( f(6k) + 6 f(3k) + 8 f(-3k) + 2 \Re \{ -4(1+2\sigma) f(\rho_2 k) + 2(1-\sigma) f(\rho_5 k) - (2+7\sigma) f(2 \rho_2 k)\}   \right) \ge 0,
\end{equation}

\begin{equation}
A_3=\frac{20}{81} \left( f(6k) - 3 f(3k) - f(-3k) + 2\Re \{ -(7+2\sigma) f(\rho_2 k) + (2+\sigma) f(\rho_5 k) + (1+2\sigma) f(2 \rho_2 k)\}   \right) \le 0,
\end{equation}

\begin{equation}
A_4=\frac{20}{81} \left( f(6k) - 3 f(3k) - f(-3k) + 2\Re \{ (5+7\sigma) f(\rho_2 k) - (1+2\sigma) f(\rho_5 k) - (2+\sigma) f(2\rho_2 k)  \}   \right) \le 0,
\end{equation}

\begin{equation}
A_5=\frac{2}{81} \left( f(6k) + 15 f(3k) - 10 f(-3k) + 2\Re \{ -10(1+2\sigma) f(\rho_2 k) - (4+5\sigma) f(\rho_5 k) - 5 (1-\sigma) f(2 \rho_2 k) \} \right) \le 0,
\end{equation}

\begin{equation}
A_6=\frac{5}{81} \left( f(6k) +6 f(3k) +8 f(-3k) + 2\Re \{ (8+4\sigma) f(\rho_2 k) + (2+4\sigma) f(\rho_5 k) + (2\sigma-5) f(2 \rho_2 k) \} \right) \ge 0,
\end{equation}

\begin{equation}
A_7=\frac{20}{81} \left( f(6k) -3 f(3k) - f(-3k) + 2\Re \{ (2-5\sigma) f(\rho_2 k) +(\sigma-1) f(\rho_5 k) + (1-\sigma) f(2 \rho_2 k) \} \right) \ge 0, \label{A7d4}
\end{equation}

\begin{equation}
A_8=\frac{5}{81} \left( f(6k) +6 f(3k) +8 f(-3k) + 2\Re \{ 4(\sigma-1) f(\rho_2 k) -2(2+\sigma) f(\rho_5 k) +(7+5\sigma) f(2 \rho_2 k) \} \right) \ge 0,
\end{equation}

\begin{equation}
A_9=\frac{2}{81} \left( f(6k) +15 f(3k) -10 f(-3k) + 2\Re \{ 10(\sigma-1) f(\rho_2 k) +(4\sigma-1) f(\rho_5 k) -5(1+2\sigma) f(2 \rho_2 k) \} \right) \le 0. \label{A9d4}
\end{equation}

All inequalities, numerically checked, hold for $0 \le k \le 0.616604$, where $f(-3k)$ is analytical, except Eq. (\ref{A1d4}), that holds for $0 \le k \le 0.540037$, and Eq. (\ref{A7d4}), that holds for $0 \le k \le 0.558283$. Again, above these points we get $A_1 < 0$ and $A_7 < 0$ respectively. As we shall show, the upper bound for the coupling constant lies in the domain of validity of Eqs. (\ref{A1d4})-(\ref{A9d4}).

\section{Application of the correlation inequalities}

In order to derive the coupling bounds for the area decay we use the same procedure as in the mean field derivation (section 2, equation (\ref{eq4})). Applying the same procedure to equations (\ref{SD3d}) and (\ref{SD4d}), we will prove the following statements.

Let $\beta $ be such that,

\vspace{0.5cm}
(a) $\textrm{  } 4(A_1 + A_2 + A_5) < 1$. Then, for $d = 3$
\begin{equation}
\langle W( C )\rangle  \leq  e^{-  \mid \ln( 4(A_1 + A_2 + A_5) \mid A} \label{eq29}.
\end{equation}

We numerically evaluate the condition $4(A_1 + A_2 + A_5) < 1$, resulting in $k < 0.440761$. It is worth to notice that inequalities (\ref{A1d3})-(\ref{A5d3}) are valid within the interval $0 \leq k \leq 0.440761$.

\vspace{0.5cm}
(b)$\textrm{  }6(A_1 + A_2 + A_6 + A_7 + A_8) < 1$. Then, for $d = 4$
\begin{equation}
\langle W( C )\rangle  \leq  e^{-  \mid \ln( 6(A_1 + A_2 + A_6 + A_7 + A_8) \mid A} \label{eq30}.
\end{equation}
Numerically evaluating the condition $6(A_1 + A_2 + A_6 + A_7 + A_8) < 1$ we obtain $k < 0.101220$, the upper bound for the area decay. All inequalities expressed by Eqs. (\ref{A1d4})-(\ref{A9d4}) hold in the domain $0 \le k \le 0.101220$.

\vspace{0.5cm}
Let us prove result (a). For b $ \epsilon$  C,
\begin{eqnarray}
\langle W(C)\rangle &=&
A_1 \sum_{i} \langle W(C) \chi^2_i \rangle 
+ A_2 \sum_{i<j} \langle W(C) \chi_i \chi_j \rangle 
+ A_3 \sum_{i<j<k} \langle W(C) \chi_i \chi^2_j \chi^2_k \rangle \nonumber \\
&+& A_4 \sum_{i<j<k<l} \langle W(C) \chi_i \chi_j \chi_k \chi^2_l \rangle  
+ A_5 \sum_{i<j<k<l}\langle W(C) \chi^2_i \chi^2_j \chi^2_k \chi^2_l\rangle \nonumber\\
&&\leq A_1 \sum_{i} \langle W(C) \chi^2_i \rangle 
+ A_2 \sum_{i<j} \langle W(C) \chi_i \chi_j \rangle 
+ A_5 \sum_{i<j<k<l}\langle W(C) \chi^2_i \chi^2_j \chi^2_k \chi^2_l\rangle
\label{eq27}
\end{eqnarray}

In the last inequality we have considered the fact that $A_3$ and $A_4$ are negatives, $A_1$, $A_2$, and $A_5$ are positives, and $\langle W(C)\chi_{P_i}\chi_{P_j}\chi_{P_k} \rangle $ is positive by Griffith's first inequality(\cite{Ganikhodjaev2007,Ganikhodjaev2010}). For the other functions we make use of the inequalities which are based on $\langle S_A^2 S_B\rangle   \leq  \langle S_B\rangle $, where $S_A^2 = \prod S_i^2$ and $S_B = \prod S_i$ (\cite{Ganikhodjaev2007,Ganikhodjaev2010,Braga1993,Braga1994}).   At each application of the equality we pass to an inequality by dropping the terms; after A steps we arrive at
\begin{equation}
\langle W( C )\rangle  \leq  (A_1 + A_2 + A_5)^A \text{(sum of } 4^A \text{ terms)} \label{eq16}        
\end{equation}  
where each term is less than one.

To prove result (b) for d=4 we proceed as previously for the result (a) for d=3, using equation (\ref{SD4d}). We can drop the terms in favor of an inequality. At each stage we have six terms from each $A_1$, $A_2$, $A_6$, $A_7$, and $A_8$ terms.

\section{Final Comments}
We have presented the application of the three-state generalized Ising  variables to the lattice gauge system, where the variables occupy the bonds of the lattice and are coupled by a four-spin interaction (a plaquette).  We have presented the derivation of correlation identities, which are exact in all dimensions and we made use of correlation inequalities for the generalized Potts model to obtain the upper bounds for the area decay of Wilson's loop. The coupling constants obtained for those bounds or behaviors are calculated  for d=3 and d=4. We obtain rigorous results that improve mean field calculations.

\vspace{1cm}

\end{document}